The Tarpon Corp has $200,000 of debt outstanding, and it pays an interest rate of 8% annually. Its annual sales are $800,000, its average tax rate is 25%, and its net profit margin on sales is 10%. If the company does not maintain a times interest earned (TIE) ratio of at least 5 to 1, then its bank will refuse to renew the loan and bankruptcy will result. Holding sales constant, at what operating (EBIT) margin would the bank refuse to renew the loan? Show your answer in this format: 12.34%
Answer #1

TIE = EBIT/Interest

Interest Expense = 200,000*8% = $16,000

Minimum required TIE = 5:1

Hence, Minimum EBIT Required = 16,000*5 = $80,000

OPerating profit margin = Operating Profit/Sales

= 80,000/800,000

= 10%

Hence, bank will refuse loan below operating margin of 10%